5th December 2014
The machine is nearly fully cold. The powering tests of the magnet circuits are in full swing. The big issue here is the need to train the superconducting dipoles to the 6.5 TeV current level. Training involves ramping up the dipoles in a given sector until a quench (transition to the resistive state) of a single dipole is detected. Protection systems swing into action, energy is extracted from the circuit, the current ramps down, and we go around the loop again. Next time round the quenched magnet now holds the current (i.e. it has trained), we go to a higher current and another one of the 154 dipoles in the circuit quenches... 11080 kA is the target current (6.5 TeV with a bit of margin).
Progress is nicely illustrated here. Or below (from 11th December) if you can't access the site.
4th July 2014: The splice consolidation campaign has finished after a remarkable effort from all involved. Sector 67 is at 20 K awaiting electrical continuity tests. Sector 81 is in cool-down (average temperture 210 K). Pressure tests, power converter tests, access system tests etc. ongoing elsewhere in the ring.
Linac2, Booster and PS have restarted with beam.
The first 6 weeks of 2013 were devoted to a proton-lead run. Four weeks of magnet tests led into long shutdown 1 (LS1). During LS1 there will be consolidation and upgrades of all major systems. One principal activity is the opening of all magnet interconnects, the measurement of the resistance of all splices in the cables that carry the current between the main dipole and quadrupole magnets, and the repair of any non-conformities.. There are around 10,000 splices in all.
Recommissioning with beam is foreseen for the start of 2015.
During operations with beam, there is a meeting every morning (8:30 during the week, 9:00 Saturday and Sunday). A record of these meetings can be found below.